Quantitative Problem: You are given the following information for Wine and Cork Enterprises (WCE):

rRF = 2%; rM = 10%; RPM = 8%, and beta = 1

What is WCE's required rate of return? Round your answer to 2 decimal places. Do not round intermediate calculations.
%

If inflation increases by 3% but there is no change in investors' risk aversion, what is WCE's required rate of return now? Round your answer to two decimal places. Do not round intermediate calculations.
%

Assume now that there is no change in inflation, but risk aversion increases by 1%. What is WCE's required rate of return now? Round your answer to two decimal places. Do not round intermediate calculations.
%

If inflation increases by 3% and risk aversion increases by 1%, what is WCE's required rate of return now? Round your answer to two decimal places. Do not round intermediate calculations.
%

EXPECTED AND REQUIRED RATES OF RETURN

Assume that the risk-free rate is 4.5% and the market risk premium is 4%.

  1. What is the required return for the overall stock market? Round your answer to two decimal places.
    %

  2. What is the required rate of return on a stock with a beta of 0.6? Round your answer to two decimal places.
    %
Answer #1

REquired rate of return = Rf + Beta ( Rm - Rf)

where Rf = risk free rate , Rm - Rf = market risk premium

Given data,

RF = 2%; rM = 10%; RPM = 8%, and beta = 1

Required rate of return= 2+1*(10-2) = 10%

----------------------------

When inflation increases by 3%,

new risk free (Rf) = 2% + 3% = 5%

Required rate of return = 5 + 1(10-5)

= 10%

----------------------------------

when risk aversion increases by 1% , beta will reduce by 1% as beta value represents risk. Lower risk means lower beta

new beta = 1-0.01 = 0.99

required rate of return = 2 + 0.99(10-2)

=2+7.92

=9.92%

--------------------------------------------

when inflation increases by 3% and risk aversion increases by 1%,

risk-free = 5% , BETA =0.99

required rate of return = 5+0.99(10-5)

= 9.950%

Hope this helps!!